Please help me!!
Katie kayaks 20 km upstream in 3 hours and returns in 1.5 hours. How fast is the stream flowing and how fast does Katie paddle in still water?

Answer by robertb(5830)

Let b = speed of the boat in still water, and s = speed of the stream.
Then 3(b-s) = 20 for upstream travel, and
and 1.5(b+s) = 20 for downstream travel.
==> b - s = 20/3 and b+s = 40/3, or
b = 10 kph.
b + s = 40/3 ==> s = 10/3 kph.